# ExoMiner++ 2.0: Vetting TESS Full-Frame Image Transit Signals

### v1

@@ VI Vetting FFI TCEs
Figure 9 shows the distribution of the orbital period against planet radius for TCEs whose scores were above 0.5. As expected, most TCEs have an orbital period shorter than the duration of a single sector. This occurs because the FFI dataset currently lacks multi-sector runs and yields on-average lower SNR for longer-period TCEs due to the fewer observations available which in turn lowers the model’s predicted scores for those examples. The longest-period TCE ($\approx 106$ days) has a mean score of about 0.59 and an estimated planet radius of approximately $11.5\ R_{\oplus}$. A cluster of TCEs with scores greater than 0.9 is visible for orbital periods in the range of $[2,5]$ days and planet radius in the range of $10,20]\ R_{\oplus}$, consistent with the expected parameters for hot Jupiters. For the 58 TCEs in the range of potential terrestrial planets ($R_{\oplus}<1.7$), there is no clear pattern in their distribution of period and planet radius.
@@
VII Conclusions
In this work, we extended the ExoMiner++ framework to vet TCEs detected in TESS FFI data. By adapting the original architecture and preprocessing pipeline to account for the unique characteristics of FFI observations (such as longer cadence and increased flux contamination), we demonstrated that deep learning models can generalize effectively beyond the 2-minute cadence domain. Key improvements included the integration of neighboring star information, higher-resolution difference images, and additional image channels such as difference image SNR, which collectively enhanced the model’s ability to identify off-target transit sources. Our experiments using multi-source learning across both 2-minute and FFI datasets showed that leveraging the larger and better-labeled 2-minute dataset improves performance on FFI data, mitigating the limitations imposed by fewer multi-sector runs and noisier light curves. The new architecture achieved reasonable precision and recall under stringent operating points, enabling reliable discrimination between planets and false positives. Although the model is more conservative than its predecessor by favoring precision over recall, it reduces misclassification of nearby eclipsing binaries and other astrophysical false positives. Importantly, it appropriately lowers confidence in crowded fields where bright neighboring stars are present, reflecting a desirable caution in ambiguous scenarios.
Applying the trained ensemble to unlabeled FFI TCEs produced a ranked catalog of candidates with associated uncertainty estimates, offering a valuable resource for prioritizing follow-up observations. This catalog expands the pool of vetted candidates beyond the 2-minute targets, supporting population studies and enabling the discovery of planets in regions of parameter space previously underexplored.
Future work will focus on optimizing the difference image branch through dedicated optimization and attention mechanisms, refining hyperparameters, and incorporating future sector data. These enhancements, together with the model’s ability to scale to large datasets, will help ensure that ExoMiner++ remains a practical and effective tool for automated vetting in current and future transit surveys.
@@ Acknowledgments
MM and HV are supported through TESS XRP 2022 contract 22-XRP22_2-0173, NASA Academic Mission Services (NAMS) contract number NNA16BD14C as well as the Intelligent Systems Research and Development-3 (ISRDS-3) Contract 80ARC020D00100. DC, JT, MJ, and BT are supported through NASA Cooperative Agreement 80NSSC21M0079.
@@
  • All convolutional and FC layers are followed by a parametric rectified linear unit activation (Fukushima, 1969, pReLU).
@@
  • All layer normalization layers use $\epsilon=1\mathrm{e}{-3}$, with center and scale parameters initialized to ‘zeros’ and ‘ones’, respectively.

### v2

@@ VI Vetting FFI TCEs
Figure 9 shows the distribution of the orbital period against planet radius for TCEs whose scores were above 0.5. As expected, most TCEs have an orbital period shorter than the duration of a single sector. This occurs because the FFI dataset currently lacks multisector runs and yields on-average lower SNR for longer-period TCEs (i.e., period $\gtrsim 27$ days) due to the fewer observations available which in turn lowers the model’s predicted scores for those examples. The longest-period TCE ($\approx 106$ days) has a mean score of about 0.59 and an estimated planet radius of approximately $11.5\ R_{\oplus}$. A cluster of TCEs with scores greater than 0.9 is visible for orbital periods in the range of $[2,5]$ days and planet radius in the range of $10,20]\ R_{\oplus}$, consistent with the expected parameters for hot Jupiters. For the 58 TCEs in the range of potential terrestrial planets ($R_{\oplus}<1.7$), there is no clear pattern in their distribution of period and planet radius.
@@
VII Limitations
ExoMiner++ 2.0 demonstrates competent vetting capability on FFI data, but it remains subject to several practical limitations. First, the 21″ TESS pixel scale limits the ability to distinguish on‑target from nearby contaminating sources, even with increased image resolution for the encoding of neighboring stars. Second, FFI light curves exhibit greater variability and fewer multisector runs than 2‑minute data, reducing diagnostic richness and contributing to lower overall performance on FFI TCEs. Third, the labeled FFI set remains relatively small, particularly for rare subclasses, limiting the robustness of subclass‑specific recall estimates. Finally, although the architecture introduces multiple enhancements, it has not undergone a dedicated hyperparameter optimization, and the ‘Difference Image’ branch likely has room for further refinement.
VIII Conclusions
In this work, we extended the ExoMiner++ framework to vet transit signals detected in TESS FFI data, introducing architectural, preprocessing, and data‑integration improvements tailored to the unique characteristics of more contamination‑prone FFI observations. By leveraging multi‑source learning, incorporating neighboring star information, and redesigning model components, ExoMiner++ 2.0 achieves effective planet versus non‑planet discrimination across both FFI and 2‑minute datasets. The resulting vetted catalog of unlabeled FFI TCEs expands the accessible parameter space for TESS transit searches and provides a valuable resource for prioritizing follow‑up observations. Key contributions of this work include:
  • First incorporation of neighboring‑star information into a TESS machine‑learning vetter. We introduce an image‑based representation encoding the relative brightness and location of known TIC neighbors, enabling the model to reason about off‑target transit sources.
  • Architectural improvements to ExoMiner++. We merge transit-view branches, remove ‘Momentum Dump’ branch, add global average pooling for parameter efficiency, introduce learned downsampling, incorporate batch and layer normalization, and strengthen residual connections for more stable training.
  • Multi‑source learning across 2‑minute and FFI datasets. The model jointly learns from both domains, improving FFI performance by leveraging the richer labels and cleaner systematics of the 2‑minute dataset.
  • A vetted catalog of 150k previously unlabeled FFI TCEs. We generate mean predictions and uncertainty estimates from an ensemble of CV models, offering a ranked list of candidates for community follow‑up, population studies, and TOI augmentation.
  • Improved sensitivity to off‑target and contamination‑driven false positives. ExoMiner++ 2.0 systematically reduces the scores of transit‑like signals originating near bright neighbors, making it more conservative in crowded fields.
Together, these advancements establish ExoMiner++ 2.0 as a robust, scalable, and contamination‑aware vetting system for TESS FFI transit searches. As additional FFI sectors become available, we expect further improvements through expanded multi‑source training, dedicated hyperparameter optimization of the ‘Difference Image’ branch, and the incorporation of attention mechanisms.
@@ Acknowledgments
MM and HV are supported through TESS XRP 2022 contract 22-XRP22_2-0173, NASA Academic Mission Services (NAMS) contract number NNA16BD14C as well as the Intelligent Systems Research and Development Support-3 (ISRDS-3) Contract 80ARC020D00100. DC, JT, MJ, and BT are supported through NASA Cooperative Agreement 80NSSC21M0079.
@@
  • All convolutional and FC layers are followed by a parametric rectified linear unit activation (He et al., 2015, pReLU).
@@
  • Layer normalization layers were applied before the classification head and after each FC layer in the classification head. All layer normalization layers use $\epsilon=1\mathrm{e}{-3}$, with center and scale parameters initialized to ‘zeros’ and ‘ones’, respectively.